# Carlo Orientale Caputo SISSA - International School for Advanced Studies, 34136 Trieste, Italy Elias Seiffert

### v3

@@ Appendix B Existence and uniqueness of the solution for $\alpha$
Interestingly, a solution only exists for $\xi<1$, i.e. for $g>g_{c}$, and in the limit $g\to g_{c}$ the fixed point distribution tends to $u_{n}$. Eq. (38) has the same form of an HFM with $m>n$ features, marginalised over the $m-n$ most detailed ones (see Eq. 5). The value of $m$ can be computed equating $1-\beta$ to the coefficient of $u_{n}$ in the marginal of $h_{m}(\mathbf{s}_{1:m})$ over $\mathbf{s}_{1:n}$. This yields the equation
@@
Appendix D TAP solutions
The Thouless-Anderson-Palmer (TAP) equations [46] are the local minima of the TAP free energy, that can be obtained with a high temperature expansion of the Legendre transform of the free energy of an extended system with extra fields on each spin variable. For a RBM the extended free energy takes the form [47]:
$-\beta F(\boldsymbol{\phi},\boldsymbol{\psi})=\log\sum_{\mathbf{x,s}}\exp\bigg{[}-\beta E(\mathbf{x,s;W,b,c})+\sum_{i}\phi_{i}x_{i}+\sum_{\mu}\psi_{\mu}s_{\mu}\bigg{]}\,.$ (40)
The Legendre transform rephrases the problem in the space of magnetizations:
$-\beta\Gamma_{\beta}(\mathbf{m^{x},m^{s}})=-\beta\max_{\boldsymbol{\phi},\boldsymbol{\psi}}\bigg{[}F(\boldsymbol{\phi},\boldsymbol{\psi})+\sum_{i}\phi_{i}m_{i}^{x}+\sum_{\mu}\psi_{\mu}m^{s}_{\mu}\bigg{]}$ (41)
The TAP equations for the RBM magnetizations can be obtained from the stationary conditions of a second order expansion of $\Gamma_{\beta}(\mathbf{m^{x},m^{s}})$ around $\beta=0$, as was done in [47]:
$m_{j}^{s}$ $=\sigma\bigg{[}b_{j}+\sum_{i}W_{ij}m_{i}^{x}-W_{ij}^{2}\bigg{(}m_{j}^{s}-\frac{1}{2}\bigg{)}(m_{i}^{x}-(m_{i}^{x})^{2})\bigg{]}$
$m_{i}^{x}$ $=\sigma\bigg{[}c_{i}+\sum_{j}W_{ij}m_{j}^{s}-W_{ij}^{2}\bigg{(}m_{i}^{x}-\frac{1}{2}\bigg{)}(m_{j}^{s}-(m_{j}^{s})^{2})\bigg{]}$ (42)
Solutions to these equations for a given layer can be found by using an iterative algorithm. We use as initialization the clamped activations of each layer for each datapoint in the MNIST datatset.

### v4

@@ B.1 Existence and uniqueness of the solution for $\alpha$
Interestingly, a solution only exists for $\xi<1$, i.e. for $g>g_{c}$, and in the limit $g\to g_{c}$ the fixed point distribution tends to $u_{n}$. Eq. (41) has the same form of an HFM with $m>n$ features, marginalised over the mnm-n most detailed ones (see Eq. 24). The value of $m$ can be computed equating $1-\beta$ to the coefficient of $u_{n}$ in the marginal of $h_{m}(\mathbf{s}_{1:m})$ over $\mathbf{s}_{1:n}$. This yields the equation
@@
Appendix C Data and deep neural networks
This Section describes first the data used in this study and then the architectures that have been trained on them.
C.1 Data
We used four standard image datasets: MNIST [67], EMNIST letters [68], Fashion-MNIST [69], and CIFAR-10 [70] downloaded via the torchvision library [71]. Every dataset was binarized before training DBNs (not for auto-encoders) with the following procedure: For MNIST, EMNIST Letters, and FMNIST (all 28×28 grayscale), we thresholded pixel intensities at 0.5 to map inputs to {0,1}. For CIFAR-10 (originally 32×32 RGB), each image was first converted to grayscale and then binarized with the same threshold then reduced to 28×28 via a centered crop that removes a 2-pixel border on each side.
Below we report each dataset’s sizes and concise descriptions:
MNIST: 60 000 handwritten digits (10 classes, 28×28);
EMNIST Letters: 124 800 handwritten letters (26 classes, 28×28; upper/lowercase merged);
Fashion-MNIST: 60 000 clothing items (10 classes, 28×28);
CIFAR-10: 50 000 natural images (10 classes, originally 32×32 RGB).
For sequential DBN training, we constructed datasets of increasing breadth as follows: we first partitioned MNIST into six label sets containing {1, 2, 4, 6, 8, 10} digit classes; we then augmented with EMNIST Letters in three steps: (i) letters a–i, (ii) letters l–r, and (iii) the remaining letters, followed by Fashion-MNIST and, finally, CIFAR-10.
C.2 DBNs and training procedure
Figure 6: A three layer Deep Belief Network
In order to generate samples from the trained DBN we consider the connections between the top two layers as undirected, whereas all lower layers are connected to the upper layer by directed connections. This means that, in order to obtain a sample from a DBN we initialized a random configuration $\hat{\mathbf{s}}^{\ell}_{0}$, and performed alternating Gibbs sampling for $10^{6}$ steps to ensure convergence to the equilibrium distribution of the top RBM $p_{L}(\textbf{s}^{(L)},\textbf{s}^{(L-1)})$. Then we use this data to sample the states of lower layers using the conditional distribution $p(\mathbf{s}_{\ell-1}|\mathbf{s}_{\ell})$. In this way, we propagate the signal till the visible layer.
The DBN used in our experiment is the same as that used in Ref. [10]: it has a visible layer with $784$ nodes and $L=10$ hidden layers with the following number of nodes: $n_{\ell}=500,250,120,60,30,25,20,15,10$ and $5$, for $\ell=1,\ldots,10$.
We train the DBN one layer at a time, following Hinton’s prescription [44]. First, the bottom RBM is fitted to the data; the dataset $\hat{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ is then propagated to the first hidden layer to obtain samples $\hat{\mathbf{s}}^{(1)}$, which become the training set for the next RBM. This type of training procedure was proven [44] to increase a variational lower bound for the log-likelihood.
For each RBM, parameters are learned by stochastic gradient ascent on the log-likelihood, using Persistent Contrastive Divergence with $k=10$ (PCD-10), a learning rate of $0.01$, and mini-batches of size $100$ (see [72]), for $\sim 10^{5}$ epochs. With these settings the training of the DBNs does not exhibit mode collapse, which typically occurs when the dataset is strongly clustered and the persistent Markov chain fails to properly mix across all clusters.
We trained the DBN sequentially on datasets of progressively increasing breadth (see Data C.1). At each stage, we restarted training with the same settings (PCD-10, learning rate, mini-batch size, number of epochs), initializing all weights with those learned at the previous stage; we then trained the entire DBN on the extended dataset.
Decelle et al. [52] [73] showed that an RBM trained with CD-10 does not reproduce the equilibrium distribution, yet it can still serve as a good generative model when sampled out of equilibrium. Instead, persistent contrastive divergence (PCD-10), which we use, tends to learn a better equilibrium distribution. [^24] In Contrastive Divergence-k (CD-k), the Markov chain used to sample the distribution is initialized on the batch used to compute the gradient and $k$ Monte Carlo steps are performed. In Persistent Contrastive Divergence-k (PCD-k) the MCMC is initialized in the configuration of the previous epoch.
C.3 Autoencoder Architecture and Training Procedure
The autoencoders employed in this study were trained on the MNIST dataset and its variants [74]. Consequently, all models described below accept an input vector of dimension $784$.
C.3.1 Architecture
Each model implements a fully connected, feed-forward autoencoder composed of two principal components: an encoder that maps the high-dimensional input into a compact latent representation, and a decoder that attempts to reconstruct the original input from that representation. The encoder and decoder are constructed symmetrically around a central bottleneck (latent) layer, so that the decoder mirrors the encoder’s layer sizes and activation choices in reverse (see Fig. 5 a). This symmetric design facilitates interpretation of the learned mapping between input space and latent space and is a common architectural choice for classical autoencoders [45].
The number of hidden layers explored across experiments ranges from one to seven. Let $n_{0}=784$ denote the input dimension and let $n_{i}$ denote the number of neurons in the $i$-th hidden layer of the encoder. Successive encoder layers are defined by a fixed geometric reduction factor of $0.6$; specifically,
$n_{i+1}=\big\lceil 0.6\,n_{i}\big\rceil,$
where $\lceil\cdot\rceil$ denotes the ceiling function to ensure that layer sizes are integer-valued. This progressive compression produces a sequence of hidden sizes that decreases smoothly from the input dimension down to the pre-specified bottleneck dimensionality and provides a controlled way to vary model capacity and depth while keeping the reduction schedule consistent across models.
Activation functions were assigned as follows: all intermediate hidden layers employ the rectified linear unit (ReLU) nonlinearity, defined as $\mathrm{ReLU}(x)=\max(0,x)$; ReLU units are widely used owing to their favorable optimization properties (reduced saturation and mitigated vanishing gradients) [75]. Both the layer immediately preceding the bottleneck and the last layer of the decoder use a sigmoid activation $\sigma(x)=(1+e^{-x})^{-1}$. Attempts of using as output layer the linear function were made, but the networks showed a preference for memorizing datapoints instead of detecting features. Moreover, the Mean Squared Error (see C.3.2) is sensibly lower and better behaved in the case of the sigmoid function as output (see Fig. 7). The latent (bottleneck) dimension itself was treated as a variable: distinct networks were trained with latent dimensions $d=10$ and $12$ in order to study the effect of representational capacity on reconstruction quality and on the geometry of learned codes.
The decoder is symmetric to the encoder: starting from the latent layer, successive layers increase dimensionality according to the inverse schedule (mirroring the reduction factor of the encoder) until the reconstructed output of dimension $784$ is produced.
A symmetric architecture reduces the number of arbitrary design choices and often yields decoder features that are naturally aligned with encoder features, easing interpretability of the coding/decoding mapping. In classical autoencoders one can choose to tie decoder weights to the transpose of encoder weights (reducing parameter count and imposing linear symmetry) or allow untied weights (more flexibility at the cost of more parameters). The present work uses untied weights, which provides additional decoder capacity and avoids forcing a linear transpose constraint; the choice was motivated by the aim of assessing how representational dimensionality (bottleneck size) and depth affect reconstruction independent of a tied-weight regularization.
Figure 7: The Mean Squared Error of the autoencoders networks with ten latent dimensions, for the case of sigmoid as output layer (left) and the case of a linear layer (right). In the latter case, the increase of the error for networks with more than three hidden layers coincides with the preference of the network for memorizing data points, rather than detecting features.
C.3.2 Training procedure
All networks were trained using the Adam optimizer with learning rate $\eta=10^{-3}$ and $L_{2}$ weight decay (also referred to as weight decay) set to $10^{-5}$. Adam is an adaptive first-order optimizer that computes individual learning rates for each parameter from estimates of first and second moments of the gradients and is well suited for stochastic optimization problems that are nonstationary and noisy [76]. The training objective used for all runs was the mean squared error between the input $x$ and its reconstruction $\hat{x}$:
$\mathcal{L}_{\mathrm{MSE}}(x,\hat{x})\;=\;\frac{1}{D}\sum_{j=1}^{D}\big(x_{j}-\hat{x}_{j}\big)^{2},$
where $D=784$ is the input dimensionality. MSE is a natural choice when the goal is to minimize Euclidean reconstruction error; for image data preprocessed to lie in $[0,1]$, MSE provides a direct measure of pixelwise fidelity. Each architecture/latent dimension configuration was trained for 10 epochs through the training set.
Layer-wise pretraining and initialization strategy
To improve optimization stability and to provide favorable initializations for deeper models, a greedy layer-wise pretraining procedure was employed. The pretraining approach used here follows the pragmatic spirit of early deep representation learning strategies in which layers are learned progressively and representations learned by shallower architectures are used to initialize deeper ones [45, 77]… The precise procedure implemented is as follows:
  1. 1.
    Train a shallow autoencoder with a small number of hidden layers (for instance, one hidden layer plus a bottleneck). Training at this stage optimizes the MSE objective for the shallow architecture starting from standard random initialization.
  2. 2.
    When the shallow model has converged, retain the learned weights for the encoder part up to the deepest layer of that shallow model.
  3. 3.
    Construct a deeper model with one additional hidden layer. Initialize the weights of the new deeper model as follows: (i) copy the encoder weights from the previously trained shallow model into the corresponding positions in the deeper model; (ii) initialize the newly added layers (both encoder and matching decoder layers) with random weights (e.g., small Gaussian i.i.d. initialization); (iii) mirror the copied encoder initializations into the corresponding decoder positions if using symmetric initialization heuristics.
  4. 4.
    Train the deeper model on the same reconstruction objective until convergence.
  5. 5.
    Repeat the expansion and initialization procedure iteratively until the target depth (up to seven hidden layers in this study) is reached.
This greedy scheme yields an initialization for each deeper architecture from parameters that have already learned useful lower-level features; empirically, such initialization can reduce the likelihood of becoming trapped in poor local minima and can accelerate convergence relative to training the deepest architecture from random initialization [45, 77]..
Regularization and normalization
Batch normalization and dropout were intentionally omitted in order to maintain full transparency of the learned latent representations and to avoid introducing additional mechanisms that could obscure the relationship between architecture depth, bottleneck dimensionality, and reconstruction characteristics. Batch normalization can significantly alter the distribution of activations during training and often improves optimization speed and stability [78] ; dropout randomly zeros activations during training and acts as an implicit model ensemble/regularizer [79]. While both are valuable tools for improving generalization in many supervised and unsupervised contexts, their use would complicate a direct, controlled investigation of how depth and bottleneck size alone influence the autoencoder’s representational geometry.